OpenStudy (anonymous):

Solve each system by using elimination 2x+4y=22 2x-2y=-8 Please show work thank you

pooja195 (pooja195):

hey there welcome to openstudy! To solve we must have an opposite so neg and pos what would you multiply the first equation by to get it neg?

OpenStudy (anonymous):

Can I help her

OpenStudy (anonymous):

2x+4y=22 2x-2y=-8 okay first of all you are going to take and make this into y=mx+b do you know this formula?